Roslin Robotics stock has a volatility of 26% and a current stock price of $49 per share. Roslin pays no dividends. The​ risk-free interest is 5%. Determine the​ Black-Scholes value of a​ one-year, at-the-money call option on Roslin stock. The​ Black-Scholes value of a​ one-year, at-the-money call option on Roslin stock is

​$______(Round to the nearest​ cent.)
Answer #1
As per Black Scholes Model
Value of call option = (S)*N(d1)-N(d2)*K*e^(-r*t)
Where
S = Current price = 49
t = time to expiry = 1
K = Strike price = 49
r = Risk free rate = 5.0%
q = Dividend Yield = 0%
σ = Std dev = 26%
d1 = (ln(S/K)+(r-q+σ^2/2)*t)/(σ*t^(1/2)
d1 = (ln(49/49)+(0.05-0+0.26^2/2)*1)/(0.26*1^(1/2))
d1 = 0.322308
d2 = d1-σ*t^(1/2)
d2 =0.322308-0.26*1^(1/2)
d2 = 0.062308
N(d1) = Cumulative standard normal dist. of d1
N(d1) =0.62639
N(d2) = Cumulative standard normal dist. of d2
N(d2) =0.524841
Value of call= 49*0.62639-0.524841*49*e^(-0.05*1)
Value of call= 6.23